OpenStudy (anonymous):

What method is better to use to solve x^3-6x^2+6x=0?

OpenStudy (helder_edwin):

u can factor the x first.

OpenStudy (anonymous):

okay.

OpenStudy (anonymous):

Likely will want quadratic formula after that.

OpenStudy (anonymous):

I did that would it be 0, plus or equal to 6?

OpenStudy (helder_edwin):

\[ \large x(x^2-6x+6)=0 \]

OpenStudy (anonymous):

Could also complete the square on the quadratic: x^2-6x+6=0 x^2-6x+9=3 (x-3)^2=3 x-3=±√3 So, the three solutions would be 0, and 3±√3.

OpenStudy (anonymous):

Ohhh i see that makes sense! would you still use that method or something else to find the solution to the equatioin x^4-64=0?